OpenStudy (anonymous):

I need an explanation as to how to solve it. F(x)= {x^2, x<-5 ax + b, -5 ≤ x ≤ 4 x+12, x>4

OpenStudy (shadowfiend):

So: \[F(x) = \begin{cases}x^2 & x < -5\\ax + b & -5 \leq x \leq 4\\x + 12 & x > 4\end{cases}\]

OpenStudy (shadowfiend):

Are you trying to solve for a and b?